OpenStudy (anonymous):

Write the trigonometric expression as an algebraic expression in u. cos (cot^-1 u)

OpenStudy (anonymous):

Write the trigonometric expression as an algebraic expression in u. tan (csc-1 u)

OpenStudy (ybarrap):

$$ \cos(\arccos(u))=\cfrac{u}{\sqrt{1+u^2}} $$ This is why:  Then \(\cos \theta=\cfrac{u}{\sqrt{1+u^2}}\)

OpenStudy (ybarrap):